OpenStudy (anonymous):

WILL FAN FOR HELP <3 Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). ) (I'm sorry this is sp long and stuff im just stuck :/) (also, I know the correct answer I just cant find the way to solve it)

OpenStudy (ankitshaw):

\[3\sqrt{2x} + x \sqrt{8x} - 5\sqrt{18x} \] \[3\sqrt{2x} + 2x \sqrt{2x} - 15\sqrt{2x} \] \[\sqrt{2x}(2x-12)\] it seems that nakim took 2x-12x instead of 2x-12

OpenStudy (anonymous):

Thank you so so so so much youre a life saver <3 Actually i get it a lot more now :)

OpenStudy (ankitshaw):

welcome :)